In the commercial preparation of hydrogen chloride. what weight of HCl in grams may be obtained by heating 468 grams of NaCl with H2SO4?

2NaCl + H2SO4 = NaSO4 + 2 HCl


1
Expert's answer
2022-03-25T00:47:04-0400

Moles of NaCl heated=?

=468/58= 8.07moles

Mole ratio=1:1

Mass= 8.07× 36

= 290.48g